# Tree-Based versus Hybrid Graphical-Textual Model Editors: An Empirical Study of Testing Specifications

### v1

@@ 6.3. Preference (RQ3)
Concerning understanding a testing space that does not contain condition expressions, 12 participants reported their preference on a diagrammatic editor, 7 participants on a tree editor, and 3 participants have no preference. To explore and navigate a testing space that does not contain condition expressions, 13 participants preferred a diagrammatic editor, 7 participants a tree editor, and 2 participants have no preference. To find relationships between elements of a testing space, 19 participants preferred a diagrammatic editor, 0 participants a tree editor, and 2 participants have no preference. To understand the meaning of condition expressions, 17 participants preferred a textual representation, 3 participants a tree representation, and 2 participants have no preference. For modelling conditions from scratch, 16 participants preferred a textual representation, 3 participants a tree representation, and 3 participants have no preference. To modify existing conditions, 17 participants preferred a textual representation, 4 participants a tree representation, and 1 participant has no preference. As reported, 20 participants stated that adding a textual syntax for modelling conditions inside the tree editor would improve usability and modelling efficiency; 2 participants did not share this opinion.
The feature of the tree editor that participants particularly liked was its conciseness and hierarchical representation that evidences the order and containment of elements. Participants also noted that the tree editor provided a better sense of control over the model, enabling faster exploration, especially when using the keyboard. However, several participants reported difficulties to understand and define complex conditions with the tree editor. Many participants would have liked the tree editor to use graphical symbols and colour-coding, to be able to differentiate between element types.
@@ 6.4. Threats to Validity
Construct Validity. A construct validity threat is that participants may misinterpret some task descriptions, although they would be able to eventually understand the descriptions with the help of the session facilitator. If participants begin carrying out a task without completely understanding its description, more time would be needed to solve the task, and more mistakes would be made that may later be corrected by them. Thus, the performance metrics for the duration, correctness and errors corrected are by design influenced by the clarity and understandability of task descriptions.
Internal Validity. The main internal validity threat is that the study has been carried out by accessing a remote computer via TeamViewer. For tasks M3-M5, participants used keyboard shortcuts, e.g., control+space to display an auto-completion menu in the textual editor from the hybrid editor. Some participants faced challenges, either because TeamViewer randomly did not trigger the actions associated with the keyboard shortcuts, or because they did not know how to correctly trigger keyboard shortcuts, due to their limited knowledge of the Windows OS. In such instances, durations reported with the hybrid editor were higher than expected.
External Validity. An external validity threat is that the tree and hybrid editors have been evaluated with a single concrete syntax. Using a more appealing hybrid notation, with a different graphical representation and with a more intuitive textual syntax, may yield a better performance. Although the tasks covered a broad range of modelling activities, they were uncomplicated, as almost all participants could solve all tasks correctly. Further experiments would be useful to strengthen the generalizability of our findings.
@@ 7. Related Work
DSLs for Robotics. Model driven notations and toolchains have been used in RobotML (Dhouib et al., 2012), BRICS (Bruyninckx et al., 2013), SmartSoft (Schlegel et al., 2009), RoboChart (Miyazawa et al., 2019), and in engineering robotic perception and control  (Ciccozzi et al., 2017; Schlegel et al., 2010). Developing model-driven solutions for the robotics domain is an established area, with a common mechanism for MDE engineering using finite state machines and statecharts (Gascuena et al., 2012; Paraschos et al., 2012; Skubch et al., 2011). However, multi-robot interactions are under-studied, with (Cattivera and Casalaro, 2015) surveying studies and finding only a minority dealing with MRS rather than single robots. The FLYAQ family of graphical DSLs  (Ciccozzi et al., 2016) model the structure and behaviour of multi-robot aerial systems, considering their context (spatial and situational surroundings) and tasks assigned to robots and their behaviour. Dragule et al. (Dragule et al., 2017) extend FLYAQ with a specification language, enabling the specification of domain-specific constraints in a declarative manner and (Pinciroli and Beltrame, 2016) proposes a textual DSL for specifying the behaviour of robot swarms.
@@
Testing DSL Tools. FuzzFactory  (Padhye et al., 2019) is a DSL for developing domain-specific fuzzing applications. It generates code necessary for fuzzing applications to communicate with a fuzzer, and obtain dynamic domain-specific feedback during testing. VerifAI (Dreossi et al., 2019) proposes a complementary approach for tool-supported testing of cyber-physical systems, typically integrated with AI and machine learning approaches. It uses a runtime monitor to assess the performance of the MRS, together with a search process to automatically finds counterexamples. VerifAI uses Scenic (Fremont et al., 2019), which presents a scenario testing DSL, capable of specifying probabilistic features that can be combined to generate interesting test cases.
@@
DSL Evaluation. FQAD (Kahraman and Bilgen, 2015) introduces a methodology for DSL evaluation, considering quality characteristics like expressiveness, usability, and productivity. Using these characteristics and evaluator priorities, the process emits an assessment of DSL suitability. The technique was used in multiple case studies and found to be useful by stakeholders. However, FQAD does rely on potentially subjective user perspectives in interviews, also lacking empirical validation in textual languages. Similarly, our empirical study relies on subjective data such as user preferences; however, it is also based on objective information, e.g., correctness, number of clicks, errors corrected.
@@
DSL evaluation for multi-agent systems is used in (Challenger et al., 2016) analysing language (elements and transformations), execution and quality. The results show that 80% of final product code could be auto-generated. However, it did not consider mistakes from language users and their impact on learnability by new users. The textual DSL Athos (Hoffmann et al., 2022) for specifying transport and traffic routing simulations was evaluated to determine its effectiveness and efficiency in model creation and understanding. Athos was found useful in modelling time and test scores when used by users inexperienced in software development, and to a lesser degree in more experienced users.
@@
User comprehensibility of a business use case between textual and graphical flow chart (BPMN) representations is explored in (Ottensooser et al., 2012). All user groups could understand the textual representation, especially strong readers. Users with proxies for domain experience (business analysts) obtained increased comprehension from the graphical BPMN notation. A suggested strategy to optimise comprehension is showing both representations; first textual and then graphical. However, this may not apply directly to our study where both representations are significantly graphical, with the textual elements restricted to a subset of the syntax (condition expressions).
@@ 8. Conclusions and Future Work
Towards understanding the tradeoffs between tree-based and hybrid graphical-textual model editors, we conducted an empirical study using an industrial use case, analysing the performance, confidence and preferences of users for both model editor types. Concerning performance, there was a tradeoff between the speed of executing tasks using the editors. In particular, the hybrid editor was better for comprehending and modelling complex conditions, while the tree editor was better for analysing ordered lists of model elements. When modelling conditions, more corrected errors were recorded with the hybrid editor, however, the mistakes made were minor. Also, more keystrokes and fewer clicks were required for modelling conditions with the hybrid editor. Users were more confident with their solutions when using the hybrid editor for understanding the meaning of conditions. Finally, before the study, 50% of participants preferred a hybrid editor; at the end, the percentage increased to 73%, showing an increased preference towards the hybrid editor.

### v2

@@ 6.3. Preference (RQ3)
Concerning understanding a testing space that does not contain condition expressions, 12 participants reported their preference for a diagrammatic editor, 7 participants for a tree editor, and 3 participants had no preference. To explore and navigate a testing space that does not contain condition expressions, 13 participants preferred a diagrammatic editor, 7 participants a tree editor, and 2 participants had no preference. To find relationships between elements of a testing space, 19 participants preferred a diagrammatic editor, 0 participants a tree editor, and 2 participants had no preference. To understand the meaning of condition expressions, 17 participants preferred a textual representation, 3 participants a tree representation, and 2 participants have no preference. For modelling conditions from scratch, 16 participants preferred a textual representation, 3 participants a tree representation, and 3 participants have no preference. To modify existing conditions, 17 participants preferred a textual representation, 4 participants a tree representation, and 1 participant has no preference. As reported, 20 participants stated that adding a textual syntax for modelling conditions inside the tree editor would improve usability and modelling efficiency; 2 participants did not share this opinion.
The feature of the tree editor that participants particularly liked was its conciseness and hierarchical representation that evidences the order and containment of elements. Participants also noted that the tree editor provided a better sense of control over the model, enabling faster exploration, especially when using the keyboard. However, several participants reported difficulties in understanding and defining complex conditions with the tree editor. Many participants would have liked the tree editor to use graphical symbols and colour-coding to differentiate between element types.
@@ 6.4. Threats to Validity
Construct Validity. A construct validity threat is that participants may misinterpret task descriptions, although they could eventually understand the descriptions with the help of the session facilitator. If participants begin a task without completely understanding its description, they will need more time to solve it, and they will make more mistakes that they may later correct. Thus, the performance metrics for duration, correctness, and errors corrected are, by design, influenced by the clarity and understandability of task descriptions.
Internal Validity. The main internal validity threat is that the study was conducted by accessing a remote computer via TeamViewer. For tasks M3M5, participants used keyboard shortcuts, e.g., control+space to display an auto-completion menu in the textual editor from the hybrid editor. Some participants faced challenges, either because TeamViewer randomly did not trigger the actions associated with the keyboard shortcuts or because they did not know how to trigger keyboard shortcuts correctly due to their limited knowledge of the Windows OS. In such instances, durations reported with the hybrid editor were higher than expected.
External Validity. An external validity threat is that the tree and hybrid editors have been evaluated with a single concrete syntax. Using a more appealing hybrid notation, with a different graphical representation and with a more intuitive textual syntax, may yield a better performance. Although the tasks covered a broad range of modelling activities, they were uncomplicated, as almost all participants could solve all tasks correctly. Further experiments would be useful to strengthen the generalisability of our findings, involving hybrid editors that are not Sirius/Xtext-based and other kinds of non-EMF or custom EMF-based tree editors.
@@ 7. Related Work
DSLs for Robotics. Model driven notations and toolchains have been used in RobotML (Dhouib et al., 2012), BRICS (Bruyninckx et al., 2013), SmartSoft (Schlegel et al., 2009), RoboChart (Miyazawa et al., 2019), and in engineering robotic perception and control  (Ciccozzi et al., 2017; Schlegel et al., 2010). Developing model-driven solutions for the robotics domain is an established area, with a common mechanism for MDE engineering using finite state machines and statecharts (Gascuena et al., 2012; Paraschos et al., 2012; Skubch et al., 2011). MRS interactions rather than single robots are studied in (Cattivera and Casalaro, 2015). The FLYAQ family of graphical DSLs  (Ciccozzi et al., 2016) model the structure and behaviour of multi-robot aerial systems, considering their spatial and situational surroundings and tasks assigned to robots and their behaviour. FLYAQ is extended with a specification language in (Dragule et al., 2017), enabling the definition of domain-specific constraints in a declarative manner and (Pinciroli and Beltrame, 2016) proposes a textual DSL for specifying the behaviour of robot swarms.
@@
Testing DSL Tools. FuzzFactory  (Padhye et al., 2019) is a DSL for developing domain-specific fuzzing applications. It generates code necessary for fuzzing applications to communicate with a fuzzer, and obtain dynamic domain-specific feedback during testing. VerifAI (Dreossi et al., 2019) proposes a complementary approach for tool-supported testing of cyber-physical systems, typically integrated with AI and machine learning approaches. VerifAI uses Scenic (Fremont et al., 2019), which presents a scenario testing DSL, capable of specifying probabilistic features that can be combined to generate test cases of interest.
@@
DSL Evaluation. FQAD (Kahraman and Bilgen, 2015) introduces a methodology for DSL evaluation, considering quality characteristics like expressiveness, usability, and productivity to assess DSL suitability. Despite its usefulness in multiple case studies, FQAD relies on potentially subjective user perspectives, and also lacks empirical validation in textual languages. Similarly, our empirical study relies on subjective data such as user preferences; however, it is also based on objective information, e.g., correctness, number of clicks, errors corrected.
@@
DSL evaluation for multi-agent systems is used in (Challenger et al., 2016), analysing language (elements and transformations), execution and quality. The results show that 80% of the final product code could be auto-generated. However, it did not consider mistakes from language users and their impact on learnability by new users. The textual DSL Athos (Hoffmann et al., 2022) for specifying transport and traffic routing simulations was evaluated to determine its effectiveness and efficiency in model creation and understanding. Athos was found useful in modelling time and test scores when used by users inexperienced in software development, and to a lesser degree, in more experienced users.
@@
User comprehensibility of a business use case between textual and graphical flowchart (BPMN) representations is explored in (Ottensooser et al., 2012). All user groups could understand the textual representation (especially strong readers) and experienced domain analysts obtained increased comprehension from BPMN. This suggests comprehension can be optimised by showing both representations; first textual and then graphical. This may not apply directly to our study where both representations are significantly graphical, with the textual elements restricted to a subset of the syntax (condition expressions).
In (Jakšić et al., 2014), tests were performed to determine whether users could better create and manage feature models (FMs) using a diagrammatic representation, as compared to a textual representation. The results are aligned with our insights and showed statistically that users authored FMs of a higher quality and managed FMs with better efficiency and effectiveness with the diagrammatic representation. Finally, in (Jolak et al., 2020), comparisons were performed between a graphical UML language and textual descriptions of the associated system, a familiar MVC fitness centre architecture. It found that the graphical representation is generally better than a textual description of the associated system, which broadly corroborates the findings in our investigation.
@@ 8. Conclusions and Future Work
Our empirical study, using an industrial use case, aimed to analyse and understand users’ performance, confidence, and preferences between using tree-based and hybrid graphical-textual model editors. Concerning performance, there was a tradeoff between the speed of executing tasks using the editors. The hybrid editor was better for comprehending and modelling complex conditions, while the tree editor was better for analysing ordered lists of model elements. When modelling conditions, more corrected errors were found with the hybrid editor; however, the mistakes made were minor. Also, more keystrokes and fewer clicks were made for modelling conditions with the hybrid editor. Users were more confident with their solutions when using the hybrid editor to understand the meaning of conditions. Finally, before the study, 50% of participants preferred a hybrid editor; at the end, the percentage increased to 73%, showing an increased preference towards the hybrid editor.